OpenStudy (anonymous):

Write the equation of the line with a slope of -5 and a y- intercept of (0,3).

OpenStudy (anonymous):

use point slope formula y-y1=m(x-x1) so y-3=-5(x-0) solve for y to get it into y=mx+b form

OpenStudy (anonymous):

y -3 = -5(x+0) Or y = -5(x+0) +3

OpenStudy (anonymous):

Thank you!